bath towel formerly sold for 80.00 each are now offered as 9.00 per dozen. what is the ratio of the old price to the new price? please help.. thanks and godbless

Answer by stanbon(75887)

bath towel formerly sold for 80.00 each are now offered as 9.00 per dozen. what is the ratio of the old price to the new price? please help..
---
$9.00 per dozen = 9/12 = 3/4 = 0.75 per towel
--------------
Ratio of old price to new price = 80/(3/4) = 320:3
=====================
Cheers,
Stan H.
---------------